OpenStudy (anonymous):

#35 "A person hired a firm to build a CB radio tower. The firm charges $100 for labor for the first 10 feet. After that, the cost of the labor for each succeeding 10 feet is $25 more than the preceding 10 feet. That is, the next 10 feet will cost $125, the next 10 feet will cost $150, etc. How much will it cost to build a 90-foot tower?" I woud like to have help with this questions

OpenStudy (anonymous):

sooo, ummm. a shortcut would be ummmmm 25 times 9 plus one HunDred

OpenStudy (anonymous):

yeah thats right. i think. so its 325 bruh

OpenStudy (anonymous):

a_9 = a_1 + 8d = 100 + 8 * 25 = $300